OpenStudy (anonymous):

what is sum of first 35 terms in the series 7+9+11... a.39 b.1399 c.1645 d.1435

OpenStudy (anonymous):

d is the answer. 75 is the 35th tirm.

OpenStudy (anonymous):

term

OpenStudy (anonymous):

Here, a = 7, d = 9 - 7 = 2 So given to find the sum of 35 terms, hence n = 35 Last term (l) = a + (n - 1)d = 7 + (35 - 1)(2) = 7 + 68 = 75 So, l = 75 Use :\[Sum = n(a + l)/2\] Sum = 35(7 + 75)/2 = (35*82)/2 = 35*41 So, the sum is : 1435 is the answer...(D)